### v1

# Unveiling the Hubble Constant through Galaxy Cluster Gas Mass Fractions

Abstract
In this work, we obtain Hubble constant ($H_{0}$) estimates by using two galaxy cluster gas mass fraction measurement samples, Type Ia supernovae luminosity distances and the validity of the cosmic distance duality relation. Notably, the angular diameter distance (ADD) to each galaxy cluster in the samples is determined by combining its gas mass fraction measurement with galaxy clustering observations, more precisely, the $\Omega_{b}/\Omega_{m}$ ratio. Such a combination results in a $H_{0}$ estimate that is independent of a specific cosmological framework. In one of the samples, the gas fraction measurements were calculated in spherical shells at radii near $r_{\rm 2500}$ (44 data points), while in the other (103 data points) the measurements were obtained by integrating X-ray observations across all radii $<r_{\rm 500}$. The quantities $r_{2500}$ and $r_{\rm 500}$ define the radii within which the mean cluster density is 2500 or 500 times the critical density of the Universe at the cluster’s redshift. We find $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc at 68% CL, respectively. We also investigate the impact on the $H_{0}$ determination by exploring the precision and number of gas mass fraction data by performing a data Monte Carlo simulation. Our simulations show that future measurements could achieve a precision of up to 3.5% for $H_{0}$.
I Introduction
In contemporary cosmology, the Hubble constant, denoted as $H_{0}$, holds a pivotal position in our comprehension of the universe’s expansion and physical composition. Serving as a fundamental parameter, it delineates the rate at which the universe is currently expanding, thus playing a crucial role in theoretical frameworks aimed at deciphering its age and dynamics. However, recent years have witnessed a surge in interest and discourse concerning the precise value of the Hubble constant. Observations employing various methodologies have produced subtly divergent outcomes, resulting in tensions in the determination of $H_{0}$ Riess and Breuval (2023); Freedman and Madore (2023).
[…]
Moreover, it is crucial to measure $H_{0}$ independently of both CMB data and the local distance ladder method. In this line, one may obtain the Hubble constant via galaxy cluster observations. For instance, in Holanda et al. (2012a, 2014) it is considered galaxy cluster angular diameter distance combined with other geometrical measurements to obtain $H_{0}$ with an accuracy of 5%. Still in this context, precise measurements of $f_{\text{gas}}$ from X-ray data offer a pathway to constrain cosmological models. When combined with external data, they have yielded robust constraints on cosmological parameters. Concerning $H_{0}$, a comprehensive analysis was conducted in Holanda et al. (2020), combining galaxy cluster X-ray gas mass fraction and BAO measurements to derive precise constraints on $H_{0}$ within both the $\Lambda$CDM and $w$CDM models. Additionally, Mantz et al. (2021) presents related work in this area. Numerous other analyses utilizing galaxy clusters have been documented in the literature (see, for instance, Holanda et al. (2014, 2012b); Cunha et al. (2007); Bora and Holanda (2023); Kozmanyan et al. (2019); Barbosa et al. (2024); Panchal and Desai (2024); Darragh-Ford et al. (2023); Wicker et al. (2023)).
Recently, Colaço et al. (2023) introduced new constraints on $H_{0}$ by utilizing a combination of the Pantheon Type Ia supernova (SNIa) sample, galaxy cluster angular diameter distances, and the cosmic distance duality relation validity Etherington (1933). Their analysis involved aggregating statistical and systematic errors in galaxy cluster measurements in quadrature, resulting in $H_{0}=67.22\pm 6.07$ km/s/Mpc at the $1\sigma$ confidence level (CL). This methodology closely follows the approach previously established in Ref. Renzi and Silvestri (2023), ensuring full independence from any specific cosmological model.
Our study aims to constrain the $H_{0}$ parameter by leveraging a combination of independent datasets. We utilize unanchored luminosity distance estimates obtained from the apparent magnitude Pantheon SNIa sample along with two galaxy cluster angular diameter distance (ADD) samples. It is important to stress that the galaxy cluster ADDs are obtained by combining $f_{\text{gas}}$ measurements and galaxy clustering observations. These ADD estimates require the $\Omega_{b}/\Omega_{m}$ ratio value, so we incorporate two measurements: one from Planck Collaboration Aghanim et al. (2020) and another from galaxy clustering Krolewski and Percival (2024). It is crucial to note that the combination $f_{\text{gas}}$ plus galaxy clustering provides galaxy cluster angular diameter distances directly, being, therefore, fully independent of any specific cosmological model. As previously mentioned, our methodology closely follows that developed in Renzi and Silvestri (2023), ensuring complete independence from any particular cosmological model, only the cosmic distance duality relation validity is required. The subsequent section introduces the datasets used in our analysis and the methodology. We then proceed to discuss the statistical construction of our observables. Our findings regarding the constraints on $H_{0}$ are discussed in Section III, followed by concluding remarks in Section IV.
II Data and Methodology
As mentioned previously, the method relies on unanchored luminosity distances, denoted as $[H_{0}{D_{L}}(z)]^{\rm SN}$, and $[D_{A}(z)]^{\rm GC}$ measurements from galaxy clusters at the same redshifts. The data considered in this analysis are:
  • Gas mass fraction sample 1 (GMF1): we utilize a new sample of $f_{\text{gas}}$ measurements spanning the redshift range $0.018\leq z\leq 1.160$, compiled in Mantz et al. (2022). This dataset comprises 44 massive, hot, and morphologically relaxed galaxy clusters observed by the Chandra telescope (see Fig. 1). Notably, the selection of relaxed systems aims to minimize systematic uncertainties and scatter arising from deviations in hydrostatic equilibrium and spherical symmetry. Additionally, the gas mass fraction of these structures was derived within spherical shells at radii near $r_{2500}$ ($0.8-1.2r_{2500}$), rather than integrated across all radii ($<r_{2500}$). Here, $r_{2500}$ denotes the radius at which the mean enclosed mass density equals $2500$ times the universe’s critical density at the cluster’s redshift. This restriction is crucial for reducing uncertainties in predicting the $\gamma_{0}$ factor from hydrodynamic simulations and minimizing intrinsic scatter, thereby leading to tighter cosmological constraints (for further details, see Mantz et al. (2022)).
  • Gas mass fraction sample 2 (GMF2): We analyze a dataset consisting of 103 observations spanning a redshift range from $0.0473\geq z\geq 1.235$. This dataset comprises 12 clusters with $z<0.1$ from the X-COP survey Eckert et al. ; a set of 44 clusters within the range $0.1\geq z\geq 0.3$ Ettori et al. ; and observations at high redshifts, consisting of 47 clusters obtained by Ghirardini et al. in the range $0.4\geq z\geq 1.235$ (see Fig. 1). Here, different from the previous sample, it is important to comment that the $f_{\text{gas}}$ measurements of this sample were obtained by integrating the X-ray observations across all radii $<r_{500}$, which represents the radius of the cluster where the density of the medium exceeds 500 times the critical energy density. This dataset was curated by Corasaniti et al. .
  • Type Ia supernovae: Distance modulus measurements of Type Ia supernovae are derived from the Pantheon sample Scolnic et al. (2018). This dataset includes 1048 light curves, with determined apparent magnitudes. The observations span a redshift range from 0.01 to 2.3. In all our analyses, we consider the apparent magnitudes of the Pantheon SNe Ia sample transformed into uncalibrated luminosity distances.
The methodology for constraining the Hubble constant relies on the validity of the cosmic distance duality relation (CDDR): $D_{L}(z)=(1+z)^{2}D_{A}(z)$, where $D_{A}$ and $D_{L}$ represent the angular diameter and luminosity distances, respectively. This relation is upheld if the photon number remains conserved along null geodesics in a Riemannian space-time between the observer and the source. Over recent years, various astronomical datasets have been scrutinized to test this relation, with no significant statistical evidence supporting a violation of the CDDR Holanda et al. (2012c); Xu et al. (2022); Liao (2019); Liu et al. (2021). Consequently, the broad applicability of this relation underscores its fundamental importance in observational cosmology. Any departure from it could signal the presence of new physics or systematic errors in observations Etherington (2007); Bassett and Kunz (2004); Ellis (2007). Following Renzi and Silvestri (2023), one can write
$H_{0}=\frac{1}{(1+z)^{2}}\frac{[H_{0}D_{L}(z)]^{\rm SN}}{[D_{A}(z)]^{\rm GC}}.$ (1)
In this context, $[H_{0}D_{L}(z)]^{\rm SN}$ represents the unanchored luminosity distance (see Fig. 3). Notably, it becomes feasible to derive estimates for $H_{0}$ by utilizing measurements of unanchored luminosity distance [$H_{0}D_{L}(z)$]SN alongside angular diameter distance $[D_{A}(z)]^{GC}$ at the same redshift $z$. To achieve that, we employ Type Ia supernovae to derive [$H_{0}D_{L}(z)$]SN and leverage X-ray gas mass fraction data from galaxy clusters to determine $[D_{A}(z)]^{\rm GC}$. Further elaboration on this methodology will be provided in the following sections.
II.1 Angular diameter distance from galaxy cluster gas mass fraction
In our analyses, we focus on the gas mass fraction observed in galaxy clusters, defined as $f_{\text{gas}}=M_{\text{gas}}/M_{\text{tot}}$, where $M_{\text{gas}}$ represents the mass of intracluster gas and $M_{\text{tot}}$ denotes the total mass, encompassing both baryonic gas and dark matter. It is worth noting that baryonic matter within galaxy clusters is primarily composed of the X-ray-emitting intracluster gas Sarazin (1986). The gas mass $M_{\text{gas}}(<R)$ observed within a radius $R$ through X-ray observations can be expressed as:
$M_{gas}(<R)$ $=$ $\left(\frac{3\pi\hbar m_{e}c^{2}}{2(1+X)e^{6}}\right)^{1/2}\left(\frac{3m_{e}c^{2}}{2\pi k_{B}T_{e}}\right)^{1/4}m_{H}$ (2)
×$\mbox{}\times\frac{1}{[\overline{g_{B}}(T_{e})]^{1/2}}{r_{c}}^{3/2}\left[\frac{I_{M}(R/r_{c},\beta)}{I_{L}^{1/2}(R/r_{c},\beta)}\right][L_{X}(<R)]^{1/2}\;,$
where $m_{e}$ and $m_{H}$ are the electron and hydrogen masses, respectively, $\overline{g_{B}}(T_{e})$ is the Gaunt factor, $X$ is the hydrogen mass fraction, $T_{e}$ is the gas temperature, $r_{c}$ stands for the core radius,
[…]
and
[…]
By assuming that the intracluster gas is in hydrostatic equilibrium, the total mass within a given radius $R$ can be obtained via Allen et al. (2011a)
$M_{\text{tot}}(<R)=-\left.\frac{k_{B}T_{e}R}{G\mu m_{H}}\frac{d\ln n_{e}(r)}{d\ln r}\right|_{r=R}.$ (3)
Figure 1: Gas mass fraction data, $f_{\text{gas}}$, for the two samples considered in this work.
The expected constancy of $f_{\text{gas}}$ within massive, hot, and relaxed galaxy clusters is commonly utilized to constrain cosmological parameters. In analyses using the galaxy cluster sample from Ref. Mantz et al. (2022), cosmological studies are achieved using the following equation (referenced in, for example, Allen et al. (2008); Ettori et al. (2009); Allen et al. (2011b); Mantz et al. (2014); Holanda et al. (2019); Mantz et al. (2022)):
$f_{\text{gas}}(z)=\gamma(z)K(z)A(z)\left[\frac{\Omega_{b}}{\Omega_{m}}\right]\left(\frac{D_{A}^{*}}{D_{A}}\right)^{3/2}.$ (4)
In this expression, $D_{A}^{*}$ denotes the angular diameter distance to the galaxy cluster used in the observations to obtain $f_{\text{gas}}$ (a flat-$\Lambda$CDM model with Hubble constant $H_{0}=70$ km s-1 Mpc-1 and the present-day total matter density parameter $\Omega_{m}=0.3$), $K(z)$ stands for the mass calibration and $A(z)$ represents the angular correction factor, which is close to unity for all cosmologies and redshifts of interest, and it can be neglected without significant loss of accuracy Allen et al. (2008). The $\gamma(z)$ factor is the depletion parameter, which indicates the amount of cosmic baryons that are thermalized within the cluster potential (Allen et al. (2008); Ettori et al. (2009); Allen et al. (2011b); Mantz et al. (2014, 2022)). It is worth to comment that cosmological analyses with gas mass fraction measurements are model-independent due to the ratio in the parenthesis of Eq. (4), which takes into account the expected variation in the gas mass fraction measurement when the underlying cosmology is varied.
In analyses using the GMF2, where we followed the modeling of the $f_{\text{gas}}$ given by Allen et al. ,
$f_{\text{gas}}(z)=K\gamma\left(\frac{\Omega_{b}}{\Omega_{m}}\right)\left[\frac{D_{A}^{*}(z)}{D_{A}(z)}\right]^{3/2}-f_{*},$ (5)
where $f_{*}$ is the stellar fraction. We assume a Gaussian prior on the stellar fraction, such as $f_{*}=0.015\pm 0.005$. This value is consistent with estimates from a sample of clusters with masses in the same range as those of the gas mass fraction dataset Eckert et al. . The GMF1 and GMF2 samples are shown in Fig. 1.
From Eqs.(4) and (5) it is straightforward that
$D_{A}(z)=\left(\frac{\gamma(z)K(z)A(z)\Omega_{b}}{f_{\rm gas}(z)\Omega_{m}}\right)^{2/3}D_{A}^{*},$ (6)
and
$D_{A}(z)=\left(\frac{\gamma(z)K(z)}{(f_{\rm gas}(z)+f_{*})}\frac{\Omega_{b}}{\Omega_{m}}\right)^{2/3}D_{A}^{*}.$ (7)
Figure 2: Galaxy cluster angular diameter estimates from gas mass fraction measurements (GMF1 and GMF2) by using $\Omega_{b}/\Omega_{M}$ obtained via galaxy clustering observations Krolewski and Percival (2024).
The determination of the angular diameter distance relies on the $\Omega_{b}/\Omega_{m}$ ratio. To derive the $D_{A}(z)$ value for each galaxy cluster, we employ two approaches. Firstly, we use the most recent determination of this ratio obtained from galaxy clustering as presented in Krolewski and Percival (2024). Additionally, we perform our analyses by considering this ratio from the Planck-CMB data (TT+TE+EE+lowE+lensing) Aghanim et al. (2020). The combination of $f_{\text{gas}}$ and galaxy clustering provides a direct angular diameter distance for each galaxy cluster.
For the GMF1 sample, we consider the parameterizations $\gamma(z)=\gamma_{0}(1+\gamma_{1}z)$ and $K(z)=K_{0}(1+K_{1}z)$, with uniform priors on $\gamma_{1}$ and $K_{1}$ ranging between $-0.05$ and $0.05$ Mantz et al. (2014, 2022). Still in agreement with Mantz et al. (2014, 2022), we apply the following Gaussian priors: $\gamma_{0}=0.79\pm 0.07$ Planelles et al. (2013) and $K_{0}=0.93\pm 0.11$ Applegate et al. (2014). This mass bias was obtained by incorporating weak gravitational lensing measurements from the Weighting the Giants project Applegate et al. (2014). On the other hand, for the GMF2 sample, we use the values $\gamma(z)=0.931(1+0.017z+0.003z^{2})$ (with standard deviation $\sigma=0.04$) in full agreement with the results from the FABLE simulations Henden et al. (2020) and $K=0.84\pm 0.04$ given by the analysis of a sample of clusters in Ref. Herbonnet et al. (2020) from the Canadian Cluster Comparison Project Hoekstra et al. (2015). Naturally, the values adopted are different for each GMF sample due to the x-ray gas mass fraction measurements have been obtained by considering different galaxy cluster radii. The angular diameter estimates from these two samples are presented in Fig. 2, where we consider the previously mentioned mean values and errors of the astrophysical parameters and the $\Omega_{b}/\Omega_{M}$ obtained via galaxy clustering observations Krolewski and Percival (2024).
II.2 The unanchored luminosity distance from Pantheon SN Ia compilation
We require the unanchored luminosity distance $[H_{0}D_{L}(z)]$, which can be obtained from the apparent magnitude of SNe Ia through the following relation:
$m_{B}=5\log_{10}[H_{0}D_{L}(z)]-5a_{B},$ (8)
where $m_{B}$ is the apparent peak magnitudes of the Pantheon catalog, and we consider the intercept estimate of the Hubble diagram, $a_{B}=0.71273\pm 0.00176$ Riess et al. (2016). Such $a_{B}$ value is independent of any absolute scale of luminosity or distance, and its combination with the apparent magnitude provides distance measurements independent of the selection of the light-curve fitter, fiducial source or filter Riess et al. (2016).
To obtain a sample of uncalibrated luminosity distances, we transform the Pantheon apparent magnitude measurements with the following relation PhysRevD.107.023520:
$[H_{0}D_{L}]^{\rm SN}(z)=10^{(m_{b}+5a_{B})/5}\equiv 10^{m_{b}^{\prime}/5}.$ (9)
In order to take into account the complete systematics and statistical uncertainties of the SN data in our analysis, we add the error of the intercept of the Hubble diagram to the non-diagonal covariance matrix of the apparent magnitude,
$\textbf{Cov}(\bm{m}^{\prime}_{b})=\textbf{Cov}(\bm{m}_{b})+(5\sigma_{a{{}_{B}}})^{2}\bm{I},$ (10)
being $\bm{I}$ the unity matrix. Thus, we apply the matrix transformation relation of the $m^{\prime}_{B}$ into the unanchored luminosity distance,
$\textbf{Cov}([\bm{H}_{0}\bm{D}_{L}]^{SN})=$
$\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)\textbf{Cov}(\bm{m}^{\prime}_{b})\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)^{T},$ (11)
where $\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}$ represents the partial derivative matrix of the unanchored luminosity distance vector $[\bm{H}_{0}\bm{D}_{L}]^{SN}$ concerning the vector $\bm{m}^{\prime}_{b}$. Because the luminosity distance of a SN Ia depends only on its own apparent magnitude, this matrix is diagonal.
As it was explained, this method to estimate the Hubble constant requires measurements of $[H_{0}D_{L}]^{\rm SN}$ and $[D_{A}]^{\rm GC}$ at the same redshift. To obtain the unanchored luminosity distance at the redshift of gas mass fraction data, we perform a GP regression of the constructed $[H_{0}D_{L}]^{\rm SN}$ sample considering the covariance matrix in Eq. (11). The GP method assumes a prior mean function to describe the behavior of the data and a covariance function or kernel to quantify the correlation of the observable in different domain points. Due to its continuity and differentiability, we adopt the usual square exponential kernel given by:
$k(z,z^{\prime})=\sigma^{2}\exp\left(-\frac{(z-z^{\prime})^{2}}{2l^{2}}\right),$ (12)
where $\sigma$ and $l$ are the hyperparameters related to the variation of the estimated function and its smoothing scale, respectively. We consider the zero function as a prior mean function to reduce the bias that a cosmological model can generate. The final reconstruction is obtained from the conditional probability considering the prior mean function given the data and depends explicitly on the values of these hyperparameters. We consider the values of $\sigma$ and $l$ that maximize the marginal likelihood, or equivalently, its logarithm,
$\ln{\mathcal{L}}=-\frac{1}{2}([\bm{H_{0}D_{L}}]^{SN})^{T}[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}$
$-\frac{1}{2}\ln|\bm{K}(\bm{z},\bm{z})+\bm{C}|-\frac{n}{2}\ln 2\pi,$ (13)
where $\bm{z}$ and $[\bm{H}_{0}\bm{D}_{L}(z)]^{SN}$ are the vectors of the independent and dependent data variables, respectively, $\bm{C}$ is the covariance matrix of the data given by Eq. (11) and $n$ the number of data points. The mean value of the reconstructed function , $\overline{[\bm{H_{0}D_{L}}]}^{*}$ , given the hyperparameters obtained by maximising Eq. (13) is
$\overline{[\bm{H_{0}D_{L}}]}^{*}=\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}.$ (14)
In this case, $\bm{z}^{*}$ is the redshift vector of the $[D_{A}]^{\rm GC}$ data. These $\overline{[\bm{H_{0}D_{L}}]}^{*}$ estimates are correlated and their covariance matrix is given by:
$\bm{C}_{H_{0}D_{L}}\equiv\textbf{Cov}([\bm{H_{0}D_{L}}]^{*})=$
$\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}\bm{K}(\bm{z},\bm{z}^{*}).$ (15)
We use the GaPP [^1] https://github.com/carlosandrepaes/GaPP code to estimate the $\sigma$, $l$ and the $\overline{[\bm{H_{0}D_{L}}]}^{*}$ values Seikel et al. (2012). In Fig. 3, we present the GP reconstruction of the unanchored luminosity distance. In what follows, we present the $H_{0}$ constraints.
Figure 3: The GP reconstruction of $[H_{0}D_{L}(z)]^{\rm SN}$ from the SN Ia Pantheon data compilation.
III Results
We conduct a Bayesian statistical analysis to constrain $H_{0}$ using the Markov Chain Monte Carlo (MCMC) method. We employ the emcee sampler Foreman-Mackey et al. (2013) to generate the chains and use the GetDist Python package Lewis (2019) for chain processing.
By defining the vector $\bm{\Delta}\equiv H_{0}(1+\bm{z}^{*})^{2}[\bm{D}_{A}(\bm{z}^{*})]^{\rm GC}-[\bm{H}_{0}\bm{D}_{L}(\bm{z}^{*})]^{\rm SN}$, the $\chi^{2}$ function can be written as
$\chi^{2}=\bm{\Delta}^{T}(\bm{C}_{D_{A}}+\bm{C}_{H_{0}D_{L}})^{-1}\bm{\Delta},$ (16)
where $\bm{C}_{D_{A}}$ is the diagonal matrix of the ADD errors which elements are
$(\sigma_{D_{A},i})^{2}=\left(\frac{2}{3}\frac{H_{0}(1+z_{i})^{2}[D_{A}(z_{i})]^{\rm GC}\sigma_{f_{\rm gas},i}}{f_{{\rm gas},i}}\right)^{2}.$ (17)
Finally, we express the likelihood function of our analysis as
$\mathcal{L}(\mathrm{Data}|{\bm{\Theta}})=\frac{1}{(2\pi)^{n/2}|\bm{C}_{D_{A}}+\bm{C}_{H_{0}D_{L}}|^{1/2}}\exp\left({-\frac{1}{2}\chi^{2}}\right),$ (18)
where ${\bm{\Theta}}=\{H_{0},\Omega_{b},\Omega_{m},K_{0},K_{1},\gamma_{0},\gamma_{1}\}$ is the vector of the free parameters used in the analysis. Note that the exact set of parameters depends on the gas mass fraction sample and on the $\Omega_{b}/\Omega_{m}$ prior.
It is worth mentioning that in Eq. (18), the normalization term in the likelihood relating to the errors of the cosmological observables is not negligible due to the dependence of $\sigma_{\rm DA,i}$ on the set of parameters. Finally, the posterior distribution to be sampled in the MCMC analysis is constructed as
$P({\bm{\Theta}}|\mathrm{Data})\propto\mathcal{L}(\mathrm{Data}|{\bm{\Theta}})\times\prod_{i}\pi({\Theta_{i}}).$ (19)
where $\pi({\Theta_{i}})$ corresponds to the prior probability function considered for each parameter in Table 1.
As our primary focus is on constraining the Hubble constant, we have marginalized over all cosmological and astrophysical parameters. Notably, the expressions of $D_{A}$ in Eqs. (6) and (7) crucially depend on the ratio $\Omega_{b}/\Omega_{m}=\omega_{b}/\omega_{m}$. To account for that, we incorporate a Gaussian prior for this ratio derived from galaxy clustering constraints Krolewski and Percival (2024), alongside Gaussian priors for individual parameters provided by the Planck Collaboration Aghanim et al. (2020) (see Table 1).
Parameter Prior Observable
$H_{0}$(km/s/Mpc) $\mathcal{U}[20;120]$
$\omega_{b,0}$ $\mathcal{N}[0.02237;0.0015]$ CMB
$\omega_{m,0}$ $\mathcal{N}[0.1200;0.0012]$ CMB
$\omega_{b,0}/\omega_{m,0}$ $\mathcal{N}[0.173;0.027]$ Galaxy Clustering
Table 1: Priors used in the statistical analysis. Gaussian priors are noted with $\mathcal{N}$[mean; standard deviation], while flat priors are noted as $\mathcal{U}$[min; max].
Figure 4: Posterior distributions of $H_{0}$ (km/s/Mpc) obtained by employing $\Omega_{b}$ and $\Omega_{m}$ priors from Planck 2018 Aghanim et al. (2020), and utilizing the $\Omega_{b}/\Omega_{m}$ prior from galaxy clustering Krolewski and Percival (2024).
The results are as follows:
  • GMF1: Utilizing the galaxy clustering prior yields $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc, while employing the Planck prior results in $H_{0}=68.1^{+7.7}_{-7.0}$ km/s/Mpc.
  • GMF2: With the galaxy clustering prior, we find $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc, and with the Planck prior, $H_{0}=72.8^{+5.2}_{-5.0}$ km/s/Mpc.
The posteriors are illustrated in Fig. 4. These results underscore the significant influence of a precise determination of the ratio $\omega_{b}/\omega_{m}$ on the estimation of $H_{0}$. It is important to note that our primary findings stem from model-independent analyses utilizing galaxy cluster data. Conversely, the determination of the $\omega_{b}/\omega_{m}$ ratio from CMB experiments is contingent upon the $\Lambda$CDM model.
III.1 Estimating $H_{0}$ using simulated data
Refer to captionRefer to caption
Figure 5: Left Panel: Redshift distribution of the GMF1-like simulated data. Right Panel: Precision in $H_{0}$ [km/s/Mpc] as a function of the number of $f_{\text{gas}}$ observations. Each line represent a specific value of the average relative error of the simulated data.
In order to estimate the impact on the Hubble constant precision of the increasing number of gas mass fraction data and the improvements of its precision, we perform a Monte Carlo simulation of GMF1-like data with different relative errors and numbers of data points and apply the statistical method described above using current SN Ia data and the $\Omega_{b}/\Omega_{m}$-GC prior. To simulate the GMF1-like data, we consider the following features of the real samples:
  • The set of realizations of simulated data has the same redshift distribution as the real data, which is shown in Fig. 5.
  • We fit the redshift evolution of relative errors of the real sample taking into account the mean value and dispersion of this fit.
  • As expected that the gas mass fraction in galaxy cluster be a constant, in addition to the relative error, we also consider the dispersion of the real data around its mean value along the redshift.
Taking into account the characteristics of the real data, we perform 100 simulations with 40, 80, 160, 320, and 500 $f_{\text{gas}}$ data points. Given that the average relative error in the real data is approximately 12%, we investigate the impact of improved precision in $f_{\text{gas}}$ data by considering simulated data sets with average relative errors of 12%, 6%, and 3%, while maintaining their fitted redshift evolution. The results of this analysis are presented in the right panel of Fig. 5. Assuming that future measurements maintain the same average relative error as current samples, we can project that, with an optimistic scenario of 500 $f_{\text{gas}}$ measurements, an accuracy of 5% in $H_{0}$ could be achieved. Furthermore, if future precision improves by factors of 2 and 3 over current levels, $H_{0}$ could be measured with accuracies of 4.3% and 3.5%, respectively. A similar conclusion is found for precision in terms of GMF2-like samples. Therefore, for consistency in presentation, we have retained this analysis for our main sample, the GMF1-like data samples. Importantly, in an optimistic observational scenario, we may achieve a competitive $H_{0}$ accuracy in the near future.
IV Final Remarks
In this work, we have presented new estimates on the Hubble constant without reliance on any specific cosmological model, utilizing observations of galaxy clusters and SNe Ia. Our approach hinges solely on the validity of the cosmic distance duality relation.
For galaxy clusters, we utilized two distinct gas mass fraction measurement samples: one comprising 44 measurements (GMF1) and the other consisting of 103 measurements (GMF2), obtained from literature sources. Notably, these measurements are derived by considering different radii for these structures: $r_{2500}$ and $r_{500}$, respectively. However, to implement our methodology, we require the angular diameter distance (ADD) for each galaxy cluster. This quantity is derived by combining gas mass fraction measurements with galaxy clustering observations, specifically the $\Omega_{b}/\Omega_{m}$ ratio. This combination directly furnishes the ADD for each galaxy cluster, enabling Hubble constant estimates independent of any specific cosmological model.
Our main analysis yields the following Hubble constant estimates: $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc for GMF1 and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc for GMF2, both at 68% CL (see Fig. 4). Due to their large uncertainties, these constraints are consistent with the measurements provided by the SH0ES team and the Planck satellite, indicating potential agreement despite the inherent errors in galaxy cluster observations. It is widely recognized that a deeper understanding of intra-cluster gas and dark matter properties is essential to enhance the utility of galaxy clusters as cosmological probes.
Finally, the impact on the $H_{0}$ determination via data Monte Carlo simulation was explored. Our simulations showed that future gas mass fraction measurements could achieve a precision of up to 3.5% for $H_{0}$ (see Fig. 5). Therefore, as statistical and systematic uncertainties in galaxy cluster observations diminish, the methodology employed here, leveraging gas mass fraction measurements, holds promise for improving constraints on the Hubble constant in a minimal model-independent manner.
Acknowledgements.
RFLH thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 309132/2020-7. RCN thanks the financial support from the CNPq under project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under project No. 23/2551-0000848-3.

### v2

@@
 Abstract
-In this work, we obtain Hubble constant ($H_{0}$) estimates by using two galaxy cluster gas mass fraction measurement samples, Type Ia supernovae luminosity distances and the validity of the cosmic distance duality relation.
+In this work, we obtain Hubble constant ($H_{0}$) estimates by using two galaxy cluster gas mass fraction measurement samples, Type Ia supernovae luminosity distances, and the validity of the cosmic distance duality relation.
 Notably, the angular diameter distance (ADD) to each galaxy cluster in the samples is determined by combining its gas mass fraction measurement with galaxy clustering observations, more precisely, the $\Omega_{b}/\Omega_{m}$ ratio.
 Such a combination results in a $H_{0}$ estimate that is independent of a specific cosmological framework.
-In one of the samples, the gas fraction measurements were calculated in spherical shells at radii near $r_{\rm 2500}$ (44 data points), while in the other (103 data points) the measurements were obtained by integrating X-ray observations across all radii $<r_{\rm 500}$.
-The quantities $r_{2500}$ and $r_{\rm 500}$ define the radii within which the mean cluster density is 2500 or 500 times the critical density of the Universe at the cluster’s redshift.
-We find $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc at 68% CL, respectively.
+In one of the samples, the gas fraction measurements were calculated in spherical shells at radii near $r_{\rm 2500}$ (44 data points), while in the other (103 data points) the measurements were calculated within $r_{\rm 500}$.
+We find $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc at 68% CL for the joint analysis of these data sets.
 We also investigate the impact on the $H_{0}$ determination by exploring the precision and number of gas mass fraction data by performing a data Monte Carlo simulation.
-Our simulations show that future measurements could achieve a precision of up to 3.5% for $H_{0}$.
+Our simulations show that future measurements could achieve a precision of up to 5% for $H_{0}$.
 I Introduction
@@
 Recently, Colaço et al.
-(2023) introduced new constraints on $H_{0}$ by utilizing a combination of the Pantheon Type Ia supernova (SNIa) sample, galaxy cluster angular diameter distances, and the cosmic distance duality relation validity Etherington (1933).
+(2024) introduced new constraints on $H_{0}$ by utilizing a combination of the Pantheon Type Ia supernova (SNIa) sample, galaxy cluster angular diameter distances, and the cosmic distance duality relation validity $(1+z)^{2}D_{A}/D_{L}=1$ Etherington (1933).
 Their analysis involved aggregating statistical and systematic errors in galaxy cluster measurements in quadrature, resulting in $H_{0}=67.22\pm 6.07$ km/s/Mpc at the $1\sigma$ confidence level (CL).
@@
 II Data and Methodology
-As mentioned previously, the method relies on unanchored luminosity distances, denoted as $[H_{0}{D_{L}}(z)]^{\rm SN}$, and $[D_{A}(z)]^{\rm GC}$ measurements from galaxy clusters at the same redshifts.
+As mentioned previously, the method relies on unanchored luminosity distances, denoted as $[H_{0}{D_{L}}(z)]^{\rm SN}$, and angular diameter measurements, $[D_{A}(z)]^{\text{GC}}$, from galaxy clusters at the same redshifts.
 The data considered in this analysis are:
@@
 Here, $r_{2500}$ denotes the radius at which the mean enclosed mass density equals $2500$ times the universe’s critical density at the cluster’s redshift.
-This restriction is crucial for reducing uncertainties in predicting the $\gamma_{0}$ factor from hydrodynamic simulations and minimizing intrinsic scatter, thereby leading to tighter cosmological constraints (for further details, see Mantz et al.
-(2022)).
   • Gas mass fraction sample 2 (GMF2): We analyze a dataset consisting of 103 observations spanning a redshift range from $0.0473\geq z\geq 1.235$.
@@
 1).
-Here, different from the previous sample, it is important to comment that the $f_{\text{gas}}$ measurements of this sample were obtained by integrating the X-ray observations across all radii $<r_{500}$, which represents the radius of the cluster where the density of the medium exceeds 500 times the critical energy density.
+Here, different from the previous sample, it is important to comment that the $f_{\text{gas}}$ measurements of this sample were calculated within $r_{500}$, which represents the radius of the cluster where the density of the medium exceeds 500 times the critical energy density.
 This dataset was curated by Corasaniti et al.
@@
 In all our analyses, we consider the apparent magnitudes of the Pantheon SNe Ia sample transformed into uncalibrated luminosity distances.
-The methodology for constraining the Hubble constant relies on the validity of the cosmic distance duality relation (CDDR): $D_{L}(z)=(1+z)^{2}D_{A}(z)$, where $D_{A}$ and $D_{L}$ represent the angular diameter and luminosity distances, respectively.
+The methodology for constraining the Hubble constant relies on the validity of the cosmic distance duality relation (CDDR).
 This relation is upheld if the photon number remains conserved along null geodesics in a Riemannian space-time between the observer and the source.
@@
 Following Renzi and Silvestri (2023), one can write
-$H_{0}=\frac{1}{(1+z)^{2}}\frac{[H_{0}D_{L}(z)]^{\rm SN}}{[D_{A}(z)]^{\rm GC}}.$ (1)
-In this context, $[H_{0}D_{L}(z)]^{\rm SN}$ represents the unanchored luminosity distance (see Fig.
+$H_{0}=\frac{1}{(1+z)^{2}}\frac{[H_{0}D_{L}(z)]^{\rm SN}}{[D_{A}(z)]^{\text{GC}}}.$ (1)
+In this context, $[H_{0}D_{L}(z)]^{\text{SN}}$ represents the unanchored luminosity distance (see Fig.
 3).
-Notably, it becomes feasible to derive estimates for $H_{0}$ by utilizing measurements of unanchored luminosity distance [$H_{0}D_{L}(z)$]SN alongside angular diameter distance $[D_{A}(z)]^{GC}$ at the same redshift $z$.
-To achieve that, we employ Type Ia supernovae to derive [$H_{0}D_{L}(z)$]SN and leverage X-ray gas mass fraction data from galaxy clusters to determine $[D_{A}(z)]^{\rm GC}$.
+Notably, it becomes feasible to derive estimates for $H_{0}$ by utilizing measurements of unanchored luminosity distance [$H_{0}D_{L}(z)$]${}^{\text{SN}}$ alongside angular diameter distance $[D_{A}(z)]^{\text{GC}}$ at the same redshift $z$.
+To achieve that, we employ Type Ia supernovae to derive [$H_{0}D_{L}(z)$]${}^{\text{SN}}$ and leverage X-ray gas mass fraction data from galaxy clusters to determine $[D_{A}(z)]^{\text{GC}}$.
 Further elaboration on this methodology will be provided in the following sections.
@@
 It is worth noting that baryonic matter within galaxy clusters is primarily composed of the X-ray-emitting intracluster gas Sarazin (1986).
-The gas mass $M_{\text{gas}}(<R)$ observed within a radius $R$ through X-ray observations can be expressed as:
-$M_{gas}(<R)$ $=$ $\left(\frac{3\pi\hbar m_{e}c^{2}}{2(1+X)e^{6}}\right)^{1/2}\left(\frac{3m_{e}c^{2}}{2\pi k_{B}T_{e}}\right)^{1/4}m_{H}$ (2)
+The gas mass $M_{\text{gas}}(<R)$ observed within a radius $R$ through X-ray observations can be expressed as Sarazin (1986):
+$M_{\text{gas}}(<R)$ $=$ $\left(\frac{3\pi\hbar m_{e}c^{2}}{2(1+X)e^{6}}\right)^{1/2}\left(\frac{3m_{e}c^{2}}{2\pi k_{B}T_{e}}\right)^{1/4}m_{H}$ (2)
 ×$\mbox{}\times\frac{1}{[\overline{g_{B}}(T_{e})]^{1/2}}{r_{c}}^{3/2}\left[\frac{I_{M}(R/r_{c},\beta)}{I_{L}^{1/2}(R/r_{c},\beta)}\right][L_{X}(<R)]^{1/2}\;,$
-where $m_{e}$ and $m_{H}$ are the electron and hydrogen masses, respectively, $\overline{g_{B}}(T_{e})$ is the Gaunt factor, $X$ is the hydrogen mass fraction, $T_{e}$ is the gas temperature, $r_{c}$ stands for the core radius,
+where $m_{e}$ and $m_{H}$ are the electron and hydrogen masses, respectively, $\overline{g_{B}}(T_{e})$ is the Gaunt factor, $X$ is the hydrogen mass fraction, $T_{e}$ is the gas temperature, $r_{c}$ stands for the core radius, $L_{X}(<R)$ is the total X-ray luminosity,
 […]
@@
 […]
+This equation is derived by adopting a spherical $\beta$-model for the cluster density profile as Cavaliere and Fusco-Femiano (1978):
+$n_{e}(\theta)=n_{0}\left(1+\frac{\theta^{2}}{\theta_{c}^{2}}\right)^{-3\beta/2},$ (3)
+where $\beta$ is a dimensionless parameter that relates the kinetic energy of the galaxies within a cluster to the thermal energy of the hot gas that permeates the cluster, $\theta$ represents the radial distance from the center of the cluster, $\theta_{c}$ is the core radius and $n_{0}$ is the central electron density Sarazin (1986).
 By assuming that the intracluster gas is in hydrostatic equilibrium, the total mass within a given radius $R$ can be obtained via Allen et al. (2011a)
@@
 (2021)):
-$f_{\text{gas}}(z)=\gamma(z)K(z)A(z)\left[\frac{\Omega_{b}}{\Omega_{m}}\right]\left(\frac{D_{A}^{*}}{D_{A}}\right)^{3/2}.$ (4)
+$f_{\text{gas}}(z)=\gamma_{g}(z)K(z)A(z)\left[\frac{\Omega_{b}}{\Omega_{m}}\right]\left(\frac{D_{A}^{*}}{D_{A}}\right)^{3/2}.$ (5)
 In this expression, $D_{A}^{*}$ denotes the angular diameter distance to the galaxy cluster used in the observations to obtain $f_{\text{gas}}$ (a flat-$\Lambda$CDM model with Hubble constant $H_{0}=70$ km s-1 Mpc-1 and the present-day total matter density parameter $\Omega_{m}=0.3$), $K(z)$ stands for the mass calibration and $A(z)$ represents the angular correction factor, which is close to unity for all cosmologies and redshifts of interest, and it can be neglected without significant loss of accuracy Allen et al.
 (2008).
-The $\gamma(z)$ factor is the depletion parameter, which indicates the amount of cosmic baryons that are thermalized within the cluster potential (Allen et al.
+This quantity is independent of the Hubble constant.
+The $\gamma_{g}(z)$ factor is the gas depletion parameter, which indicates the amount of gas that is thermalized within the cluster potential (Allen et al.
 (2008); Ettori et al.
@@
 (2011b); Mantz et al.
-(2014, 2022)).
+(2014, 2021)).
+The restriction of selecting hot and relaxed clusters is crucial for reducing uncertainties in predicting the $\gamma_{g}(z)$ factor from hydrodynamic simulations and minimizing intrinsic scatter, thereby leading to tighter cosmological constraints (for further details, see Mantz et al.
+(2021)).
 It is worth to comment that cosmological analyses with gas mass fraction measurements are model-independent due to the ratio in the parenthesis of Eq.
@@
 In analyses using the GMF2, where we followed the modeling of the $f_{\text{gas}}$ given by Allen et al. ,
-$f_{\text{gas}}(z)=K\gamma\left(\frac{\Omega_{b}}{\Omega_{m}}\right)\left[\frac{D_{A}^{*}(z)}{D_{A}(z)}\right]^{3/2}-f_{*},$ (5)
+$f_{\text{gas}}(z)=K\gamma_{b}(z)\left(\frac{\Omega_{b}}{\Omega_{m}}\right)\left[\frac{D_{A}^{*}(z)}{D_{A}(z)}\right]^{3/2}-f_{*},$ (6)
 where $f_{*}$ is the stellar fraction.
 We assume a Gaussian prior on the stellar fraction, such as $f_{*}=0.015\pm 0.005$.
+The value of $\gamma_{b}(z)$ to GMF2 sample corresponds to the baryonic depletion factor and not the gas depletion.
+Therefore, it is necessary in this case to subtract the mass present in the stars in Eq.(5).
 This value is consistent with estimates from a sample of clusters with masses in the same range as those of the gas mass fraction dataset Eckert et al.
@@
 From Eqs.(5) and (6) it is straightforward that
-$D_{A}(z)=\left(\frac{\gamma(z)K(z)A(z)\Omega_{b}}{f_{\rm gas}(z)\Omega_{m}}\right)^{2/3}D_{A}^{*},$ (6)
+$D_{A}(z)=\left(\frac{\gamma_{g}(z)K(z)A(z)\Omega_{b}}{f_{\text{gas}}(z)\Omega_{m}}\right)^{2/3}D_{A}^{*},$ (7)
 and
-$D_{A}(z)=\left(\frac{\gamma(z)K(z)}{(f_{\rm gas}(z)+f_{*})}\frac{\Omega_{b}}{\Omega_{m}}\right)^{2/3}D_{A}^{*}.$ (7)
+$D_{A}(z)=\left(\frac{\gamma_{b}(z)K(z)}{(f_{\text{gas}}(z)+f_{*})}\frac{\Omega_{b}}{\Omega_{m}}\right)^{2/3}D_{A}^{*}.$ (8)
 Figure 2: Galaxy cluster angular diameter estimates from gas mass fraction measurements (GMF1 and GMF2) by using $\Omega_{b}/\Omega_{M}$ obtained via galaxy clustering observations Krolewski and Percival (2024).
@@
 The combination of $f_{\text{gas}}$ and galaxy clustering provides a direct angular diameter distance for each galaxy cluster.
-For the GMF1 sample, we consider the parameterizations $\gamma(z)=\gamma_{0}(1+\gamma_{1}z)$ and $K(z)=K_{0}(1+K_{1}z)$, with uniform priors on $\gamma_{1}$ and $K_{1}$ ranging between $-0.05$ and $0.05$ Mantz et al.
-(2014, 2022).
+For the GMF1 sample, we consider the parameterizations $\gamma_{g}(z)=\gamma_{0}(1+\gamma_{1}z)\left(\frac{M_{2500}}{3\times 10^{14}M_{\odot}}\right)^{\alpha}$ (with $\alpha=0.025\pm 0.033$) and $K(z)=K_{0}(1+K_{1}z)$, with uniform priors on $\gamma_{1}$ and $K_{1}$ ranging between $-0.05$ and $0.05$ Mantz et al.
+(2014, 2021).
 Still in agreement with Mantz et al.
-(2014, 2022), we apply the following Gaussian priors: $\gamma_{0}=0.79\pm 0.07$ Planelles et al.
+(2014, 2021), we apply the following Gaussian priors: $\gamma_{0}=0.79\pm 0.07$ Planelles et al.
 (2013) and $K_{0}=0.93\pm 0.11$ Applegate et al.
 (2014).
-This mass bias was obtained by incorporating weak gravitational lensing measurements from the Weighting the Giants project Applegate et al.
+The $\gamma_{0}$ here is from hydrodynamic simulations and corresponds directly to gas mass fraction (see Table III from Planelles et al.
+(2013)).
+The $K_{0}$ value was obtained by using 13 clusters of the sample by incorporating weak gravitational lensing measurements from the Weighting the Giants project Applegate et al.
 (2014).
-On the other hand, for the GMF2 sample, we use the values $\gamma(z)=0.931(1+0.017z+0.003z^{2})$ (with standard deviation $\sigma=0.04$) in full agreement with the results from the FABLE simulations Henden et al.
+On the other hand, for the GMF2 sample, we use the values $\gamma_{b}(z)=0.931(1+0.017z+0.003z^{2})$ (with standard deviation $\sigma=0.04$) in full agreement with the results from the FABLE simulations Henden et al.
 (2020) and $K=0.84\pm 0.04$ given by the analysis of a sample of clusters in Ref.
@@
 (2015).
-Naturally, the values adopted are different for each GMF sample due to the x-ray gas mass fraction measurements have been obtained by considering different galaxy cluster radii.
+By considering the Eq.(8) from the Ref.Angelinelli et al.
+(2022), we have multiplied this factor by $w^{\beta}x^{y+\delta*w}$, where $w=M_{500}/5\cdot 10^{14}\ h^{-1}M_{\odot}$, $x=r/R_{500,\mathrm{c}}$, while $\beta$, $\gamma$ and $\delta$ are the free parameters are $(\beta,\gamma,\delta)=(0.12,0.23,-0.22)$.
+This takes into account a possible radial and mass dependency of $\gamma_{b}(z)$.
+Naturally, the values adopted are different for each GMF sample due to the x-ray gas mass fraction measurements having been obtained by considering different galaxy cluster radii.
 The angular diameter estimates from these two samples are presented in Fig.
@@
 (2016).
-Such $a_{B}$ value is independent of any absolute scale of luminosity or distance, and its combination with the apparent magnitude provides distance measurements independent of the selection of the light-curve fitter, fiducial source or filter Riess et al.
+Such $a_{B}$ value is independent of any absolute scale of luminosity or distance, and its combination with the apparent magnitude provides distance measurements independent of the selection of the light-curve fitter, fiducial source or filter.
+This quantity has a weak dependence on cosmological models, with the estimate based on the cosmographic parameters $q_{0}$=-0.55 and $j_{0}$=1 Riess et al.
 (2016).
-To obtain a sample of uncalibrated luminosity distances, we transform the Pantheon apparent magnitude measurements with the following relation PhysRevD.107.023520:
-$[H_{0}D_{L}]^{\rm SN}(z)=10^{(m_{b}+5a_{B})/5}\equiv 10^{m_{b}^{\prime}/5}.$ (9)
+To obtain a sample of uncalibrated luminosity distances, we transform the Pantheon apparent magnitude measurements with the following relation Colaço et al. (2024):
+$[H_{0}D_{L}]^{\text{SN}}(z)=10^{(m_{b}+5a_{B})/5}\equiv 10^{m_{b}^{\prime}/5}.$ (10)
 In order to take into account the complete systematics and statistical uncertainties of the SN data in our analysis, we add the error of the intercept of the Hubble diagram to the non-diagonal covariance matrix of the apparent magnitude,
@@
 being $\bm{I}$ the unity matrix. Thus, we apply the matrix transformation relation of the $m^{\prime}_{B}$ into the unanchored luminosity distance,
-$\textbf{Cov}([\bm{H}_{0}\bm{D}_{L}]^{SN})=$
-$\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)\textbf{Cov}(\bm{m}^{\prime}_{b})\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}\right)^{T},$ (11)
-where $\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{SN}}{\partial\bm{m}^{\prime}_{b}}$ represents the partial derivative matrix of the unanchored luminosity distance vector $[\bm{H}_{0}\bm{D}_{L}]^{SN}$ concerning the vector $\bm{m}^{\prime}_{b}$.
+$\textbf{Cov}([\bm{H}_{0}\bm{D}_{L}]^{\text{SN}})=$
+$\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}\right)\textbf{Cov}(\bm{m}^{\prime}_{b})\left(\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}\right)^{T},$ (12)
+where $\frac{\partial[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}}{\partial\bm{m}^{\prime}_{b}}$ represents the partial derivative matrix of the unanchored luminosity distance vector $[\bm{H}_{0}\bm{D}_{L}]^{\text{SN}}$ concerning the vector $\bm{m}^{\prime}_{b}$.
 Because the luminosity distance of a SN Ia depends only on its own apparent magnitude, this matrix is diagonal.
-As it was explained, this method to estimate the Hubble constant requires measurements of $[H_{0}D_{L}]^{\rm SN}$ and $[D_{A}]^{\rm GC}$ at the same redshift.
-To obtain the unanchored luminosity distance at the redshift of gas mass fraction data, we perform a GP regression of the constructed $[H_{0}D_{L}]^{\rm SN}$ sample considering the covariance matrix in Eq.
+As it was explained, this method to estimate the Hubble constant requires measurements of $[H_{0}D_{L}]^{\text{SN}}$ and $[D_{A}]^{\text{GC}}$ at the same redshift.
+To obtain the unanchored luminosity distance at the redshift of gas mass fraction data, we perform a Gaussian Process (GP) regression of the constructed $[H_{0}D_{L}]^{\text{SN}}$ sample considering the covariance matrix in Eq.
 (12).
@@
 We consider the values of $\sigma$ and $l$ that maximize the marginal likelihood, or equivalently, its logarithm,
-$\ln{\mathcal{L}}=-\frac{1}{2}([\bm{H_{0}D_{L}}]^{SN})^{T}[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}$
+$\ln{\mathcal{L}}=-\frac{1}{2}([\bm{H_{0}D_{L}}]^{\text{SN}})^{T}[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{\text{SN}}$
 $-\frac{1}{2}\ln|\bm{K}(\bm{z},\bm{z})+\bm{C}|-\frac{n}{2}\ln 2\pi,$ (14)
-where $\bm{z}$ and $[\bm{H}_{0}\bm{D}_{L}(z)]^{SN}$ are the vectors of the independent and dependent data variables, respectively, $\bm{C}$ is the covariance matrix of the data given by Eq.
+where $\bm{z}$ and $[\bm{H}_{0}\bm{D}_{L}(z)]^{\text{SN}}$ are the vectors of the independent and dependent data variables, respectively, $\bm{C}$ is the covariance matrix of the data given by Eq.
 (12) and $n$ the number of data points.
@@
 (14) is
-$\overline{[\bm{H_{0}D_{L}}]}^{*}=\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{SN}.$ (14)
-In this case, $\bm{z}^{*}$ is the redshift vector of the $[D_{A}]^{\rm GC}$ data. These $\overline{[\bm{H_{0}D_{L}}]}^{*}$ estimates are correlated and their covariance matrix is given by:
+$\overline{[\bm{H_{0}D_{L}}]}^{*}=\bm{K}(\bm{z}^{*},\bm{z})[\bm{K}(\bm{z},\bm{z})+\bm{C}]^{-1}[\bm{H_{0}D_{L}}]^{\text{SN}}.$ (15)
+In this case, $\bm{z}^{*}$ is the redshift vector of the $[D_{A}]^{\text{GC}}$ data. These $\overline{[\bm{H_{0}D_{L}}]}^{*}$ estimates are correlated and their covariance matrix is given by:
 $\bm{C}_{H_{0}D_{L}}\equiv\textbf{Cov}([\bm{H_{0}D_{L}}]^{*})=$
@@
 In what follows, we present the $H_{0}$ constraints.
-Figure 3: The GP reconstruction of $[H_{0}D_{L}(z)]^{\rm SN}$ from the SN Ia Pantheon data compilation.
+Figure 3: The GP reconstruction of $[H_{0}D_{L}(z)]^{\text{SN}}$ from the SN Ia Pantheon data compilation.
 III Results
-We conduct a Bayesian statistical analysis to constrain $H_{0}$ using the Markov Chain Monte Carlo (MCMC) method.
+We conduct a joint Bayesian statistical analysis of the two aforementioned gas mass fraction data samples and the Markov Chain Monte Carlo (MCMC) method to constrain $H_{0}$ with all data available.
 We employ the emcee sampler Foreman-Mackey et al.
 (2013) to generate the chains and use the GetDist Python package Lewis (2019) for chain processing.
-By defining the vector $\bm{\Delta}\equiv H_{0}(1+\bm{z}^{*})^{2}[\bm{D}_{A}(\bm{z}^{*})]^{\rm GC}-[\bm{H}_{0}\bm{D}_{L}(\bm{z}^{*})]^{\rm SN}$, the $\chi^{2}$ function can be written as
+In order to avoid a double counting in the statistical analysis, we exclude the common clusters from the GMF2 sample: Abell2029, Abell1835, Abell2204, J1415.1+3612, RXCJ2129.6+0005, RXJ1347.5-1145 and SPT-CLJ2043-5035.
+By defining the vector $\bm{\Delta}\equiv H_{0}(1+\bm{z}^{*})^{2}[\bm{D}_{A}(\bm{z}^{*})]^{\text{GC}}-[\bm{H}_{0}\bm{D}_{L}(\bm{z}^{*})]^{\text{SN}}$, the $\chi^{2}$ function can be written as
 $\chi^{2}=\bm{\Delta}^{T}(\bm{C}_{D_{A}}+\bm{C}_{H_{0}D_{L}})^{-1}\bm{\Delta},$ (17)
 where $\bm{C}_{D_{A}}$ is the diagonal matrix of the ADD errors which elements are
-$(\sigma_{D_{A},i})^{2}=\left(\frac{2}{3}\frac{H_{0}(1+z_{i})^{2}[D_{A}(z_{i})]^{\rm GC}\sigma_{f_{\rm gas},i}}{f_{{\rm gas},i}}\right)^{2}.$ (17)
-Finally, we express the likelihood function of our analysis as
+$(\sigma_{D_{A},i})^{2}=\left(\frac{2}{3}\frac{H_{0}(1+z_{i})^{2}[D_{A}(z_{i})]^{\text{GC}}\sigma_{f_{\text{gas}},i}}{f_{{\text{gas}},i}}\right)^{2}.$ (18)
+Finally, we express the joint likelihood function of our analysis as
 $\mathcal{L}(\mathrm{Data}|{\bm{\Theta}})=\frac{1}{(2\pi)^{n/2}|\bm{C}_{D_{A}}+\bm{C}_{H_{0}D_{L}}|^{1/2}}\exp\left({-\frac{1}{2}\chi^{2}}\right),$ (19)
-where ${\bm{\Theta}}=\{H_{0},\Omega_{b},\Omega_{m},K_{0},K_{1},\gamma_{0},\gamma_{1}\}$ is the vector of the free parameters used in the analysis.
-Note that the exact set of parameters depends on the gas mass fraction sample and on the $\Omega_{b}/\Omega_{m}$ prior.
+where ${\bm{\Theta}}=\{H_{0},\omega_{b}/\omega_{m},K_{0}^{r_{2500}},K_{1}^{r_{2500}},\gamma_{0}^{r_{2500}},\gamma_{1}^{r_{2500}},\alpha,$ $K_{0}^{r_{500}},\gamma_{0}^{r_{500}}\}$ is the vector of the free parameters used in the analysis.
+It is worth to mention that we treat the gas mass calibration parameter, $K$, and the depletion factor parameter, $\gamma$, independently because each one corresponds to models with different spherical radii.
 It is worth mentioning that in Eq.
@@
 $H_{0}$(km/s/Mpc) $\mathcal{U}[20;120]$
-$\omega_{b,0}$ $\mathcal{N}[0.02237;0.0015]$ CMB
-$\omega_{m,0}$ $\mathcal{N}[0.1200;0.0012]$ CMB
+$\omega_{b,0}/\omega_{m,0}$ $\mathcal{N}[0.1564,0.001596]$ CMB
 $\omega_{b,0}/\omega_{m,0}$ $\mathcal{N}[0.173;0.027]$ Galaxy Clustering
@@
 (2020), and utilizing the $\Omega_{b}/\Omega_{m}$ prior from galaxy clustering Krolewski and Percival (2024).
-The results are as follows:
-  • GMF1: Utilizing the galaxy clustering prior yields $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc, while employing the Planck prior results in $H_{0}=68.1^{+7.7}_{-7.0}$ km/s/Mpc.
-  • GMF2: With the galaxy clustering prior, we find $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc, and with the Planck prior, $H_{0}=72.8^{+5.2}_{-5.0}$ km/s/Mpc.
+The results of the joint analysis of the GMF1 and GMF2 are as follows:
+  • Utilizing the galaxy clustering prior yields $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc.
+  • Utilizing the Planck prior, $H_{0}=79.1^{+4.3}_{-3.6}$ km/s/Mpc.
 The posteriors are illustrated in Fig.
@@
 Conversely, the determination of the $\omega_{b}/\omega_{m}$ ratio from CMB experiments is contingent upon the $\Lambda$CDM model.
+The authors of the Ref.Mantz et al.
+(2021) considered the six clusters with $z<0.16$ of the GMF1 sample, $K_{0}=0.93\pm 0.11$, a Gaussian prior on $\Omega_{b}/\Omega_{m}$ from the Planck results (considering the flat $\Lambda$CDM model) and marginalized over a non-flat cosmological model with $\omega$ free (dark energy equation parameter).
+It was estimated $H_{0}=72.22\pm 6.7$ km/s/Mpc.
+As one may see, their estimate is in full agreement with ours.
+However, no specific cosmological framework was considered in our estimates, we marginalized on a possible evolution of the astrophysical parameters or mass dependence ($\gamma(z)$ $K(z)$ and $\alpha$) and considering all galaxy clusters in a joint statistical analysis of the GMF1 and GMF2 samples.
 III.1 Estimating $H_{0}$ using simulated data
-Refer to captionRefer to caption
 Figure 5: Left Panel: Redshift distribution of the GMF1-like simulated data.
-Right Panel: Precision in $H_{0}$ [km/s/Mpc] as a function of the number of $f_{\text{gas}}$ observations.
-Each line represent a specific value of the average relative error of the simulated data.
+Right Panel: Precision in $H_{0}$ as a function of the number of $f_{\text{gas}}$ observations.
+Each line represents a specific value of the average relative error of the simulated data.
 In order to estimate the impact on the Hubble constant precision of the increasing number of gas mass fraction data and the improvements of its precision, we perform a Monte Carlo simulation of GMF1-like data with different relative errors and numbers of data points and apply the statistical method described above using current SN Ia data and the $\Omega_{b}/\Omega_{m}$-GC prior.
@@
   • We fit the redshift evolution of relative errors of the real sample taking into account the mean value and dispersion of this fit.
-  • As expected that the gas mass fraction in galaxy cluster be a constant, in addition to the relative error, we also consider the dispersion of the real data around its mean value along the redshift.
+  • As the gas mass fraction in the galaxy cluster is expected to be constant, in addition to the relative error, we also consider the dispersion of the actual data around its mean value across different redshifts.
 Taking into account the characteristics of the real data, we perform 100 simulations with 40, 80, 160, 320, and 500 $f_{\text{gas}}$ data points.
-Given that the average relative error in the real data is approximately 12%, we investigate the impact of improved precision in $f_{\text{gas}}$ data by considering simulated data sets with average relative errors of 12%, 6%, and 3%, while maintaining their fitted redshift evolution.
-The results of this analysis are presented in the right panel of Fig.
+Given that the average relative error in the real data is approximately 12%, we investigate the impact on $H_{0}$ estimates of improved precision in $f_{\text{gas}}$ data by considering simulated data sets with average relative errors of 12%, 6%, and 3%, while maintaining their fitted redshift evolution and astrophysical and cosmological parameters unaltered.
+The results of $H_{0}$ simmetrized errors Schmelling (2000) performing this analysis are presented in the right panel of Fig.
 5.
-Assuming that future measurements maintain the same average relative error as current samples, we can project that, with an optimistic scenario of 500 $f_{\text{gas}}$ measurements, an accuracy of 5% in $H_{0}$ could be achieved.
-Furthermore, if future precision improves by factors of 2 and 3 over current levels, $H_{0}$ could be measured with accuracies of 4.3% and 3.5%, respectively.
+Assuming that future measurements maintain the same average relative error as current samples, we can project that, with an optimistic scenario of 500 $f_{\text{gas}}$ measurements, an accuracy of 7.3% in $H_{0}$ could be achieved.
+Furthermore, if future $f_{\text{gas}}$ data precision improves by factors of 2 and 4 over current levels, $H_{0}$ could be measured with precision of 6.2% and 5.0%, respectively.
+It is worth mentioning that in this analysis, the precision of individual $H_{0}$ estimates obtained from each $f_{\text{gas}}$ data point is limited by the precision of the astrophysical and cosmological parameters.
+However, by combining all $H_{0}$ information statistically, the final estimate may achieve higher precisions.
 A similar conclusion is found for precision in terms of GMF2-like samples.
@@
 This combination directly furnishes the ADD for each galaxy cluster, enabling Hubble constant estimates independent of any specific cosmological model.
-Our main analysis yields the following Hubble constant estimates: $H_{0}=71.4^{+10.5}_{-8.9}$ km/s/Mpc for GMF1 and $H_{0}=75.5^{+9.9}_{-7.9}$ km/s/Mpc for GMF2, both at 68% CL (see Fig.
+Our analysis yields the following Hubble constant estimate: $H_{0}=72.7^{+6.3}_{-5.6}$ km/s/Mpc for the compilation of the samples GMF1 and GMF2 at 68% CL (see Fig.
 4).
@@
 It is widely recognized that a deeper understanding of intra-cluster gas and dark matter properties is essential to enhance the utility of galaxy clusters as cosmological probes.
+In general, the mass of a galaxy cluster obtained using strong gravitational lensing tends to be greater than the mass derived from the hypothesis of hydrostatic equilibrium.
+In other words, the hydrostatic equilibrium hypothesis adopted in these studies may be underestimating the gas mass fraction measurements and, consequently, overestimating the Hubble constant estimates.
+However, the galaxy clusters considered in analyses are limited to the most dynamically relaxed, massive clusters known.
+In turn, this restriction is critical for reducing systematic from hidrostatical equilibrium and spherical symmetry.
 Finally, the impact on the $H_{0}$ determination via data Monte Carlo simulation was explored.
-Our simulations showed that future gas mass fraction measurements could achieve a precision of up to 3.5% for $H_{0}$ (see Fig.
+Our simulations showed that future gas mass fraction measurements could achieve a precision of up to 5% for $H_{0}$ (see Fig.
 5).
@@
 Acknowledgements.
+The authors thank the referee and Susana Landau for valuable comments about this work.
 RFLH thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No.
@@
 23/2551-0000848-3.
+LRC also thanks to CNPQ support under the project No.
+169625/2023-0.
+..